### v1

@@
Does Urban Local Governance Matter? Evidence from Indiathanks: Acknowledgements: I am grateful to SHRUG, for making their data publicly available. I thank David Brasington, Gary Painter, Olivier Parent, Clemens Pilgram, and others for helpful comments and suggestions. All errors are my own. No external funding was received for this project.
Saani Rawat PhD Candidate, University of Cincinnati. Corresponding author: rawatsa@mail.uc.edu
(November 9, 2025)
This paper examines the causal effect of urban local governance on public goods provision in India. We exploit quasi-random variation in multi-threshold criteria utilized for classifying Census Towns (CTs) and focus on settlements near the thresholds that are likely to obtain statutory recognition. Using a fuzzy regression discontinuity design, we instrument for urban local governance to identify the Local Average Treatment Effect (LATE). We document a strong first stage relationship between meeting CT thresholds and statutory recognition. Our results show that obtaining an Urban Local Body (ULB) increases local public good provision: government schools increase by approximately 14 (primary), 8 (middle), and 5 (secondary), healthcare infrastructure expands by 2 hospitals and 3 family welfare centers, and financial access deepens with 15 private banks, 2 cooperative banks, and 2 agricultural credit societies. Community amenities improve modestly with an additional public library, reading room, and cinema hall. Sports infrastructure declines by 5 facilities, consistent with our understanding of reallocation of urban space and investments. Our findings suggest that timely municipalization of emerging urban areas can expand provision of certain public goods, which may improve living standards and economic opportunities in urbanizing economies.
Keywords: Urban Local Governance, Statutory towns, Census Towns, Public Goods
@@ 1 Introduction
Urban local bodies (ULBs) are responsible for a substantial share of on-the-ground public goods in developing countries—schools, primary health facilities, sanitation, streets, lighting, and local regulation. Whether and how urban local governance matters for local economic development is a long-standing question, with renewed urgency in rapidly urbanizing economies (BrolloEtAl2013; Faguet2012; PatrickMothorpe2017). A key empirical challenge is that governance structures are endogenous: places that are bigger and more developed are more likely to be recognized as urban towns and receive municipal governance, while smaller or less developed settlements may not receive such recognition and therefore stagnate even further by being outside the urban system (Gadenne2017). Simple cross-sectional comparisons in such contexts will conflate pre-existing differences with the effects of policy. In other words, urban governance status is chosen rather than randomly assigned, making causal inference difficult.
@@
This paper leverages a sharp institutional feature in India to address this identification problem: the Census Town (CT) classification thresholds that determine eligibility for an area to be considered urban. The Indian Census defines Census Towns as settlements meeting all three criteria: (i) population of at least 5,000 (ii) population density of at least 400 per square kilometer and (iii) at least 75% of the male main workforce in non-agricultural activities. Crossing these multi-dimensional cutoffs sharply increases the probability that a settlement will be formally recognized by state governments as a Statutory Town (ST) (i.e. given a municipality or other ULB status) but notably, it does not guarantee it. We exploit this quasi-random variation in recognition likelihood around the thresholds to construct a local fuzzy regression discontinuity (RD) design. In effect, meeting the Census Town criteria in 2001 serves as an instrument for obtaining urban local governance in 2011, allowing us to identify the causal effects of statutory town status on local outcomes.
@@
The core idea is a fuzzy multi-threshold RD with multiple running variables. Using rich settlement-level microdata from the Socioeconomic High-resolution Rural-Urban Geographic (SHRUG) platform (asher2019shrug), we create a running variable called frontier distance using the three CT cutoffs and restrict attention to settlements near these thresholds. We first confirm a strong first-stage relationship: becoming marginally eligible as a Census Town in 2001 leads to a large jump in the probability of statutory recognition in 2011. We then estimate the Local Average Treatment Effect (LATE) of urban local governance (statutory town status) on a broad set of development indicators measured by the 2011 Indian Census. Our outcome data are assembled by harmonizing the 2011 Village Directory (VD) and Town Directory (TD) which enumerate local infrastructure and amenities. We link these to demographic data from the 2011 Primary Census Abstract (PCA). This provides a comprehensive settlement-level panel of outcomes including educational facilities, health infrastructure, financial institutions, and other community amenities.
@@
Preview of results. We find that crossing the urban eligibility thresholds leads to a pronounced increase in the likelihood of statutory recognition. In particular, meeting all three Census Town criteria raises the probability of obtaining urban local governance by approximately 7.1 percentage points, with a first-stage F-statistic of 18.81, indicating a reasonably strong instrumental variable. Using this fuzzy RD design, our two-stage least squares (2SLS) estimates indicate that formal urban status substantially improves local public goods provision. For example, obtaining a municipality leads to an increase of about 13.5 additional government primary schools in a settlement (relative to the mean for similar settlements just below the threshold), along with roughly 7.8 more middle schools and 5.0 more secondary schools on average. We also find large positive effects on health and financial infrastructure: being recognized as a town increases the number of hospitals by about 2.5 and family welfare centers by about 3.0, and boosts the presence of banking facilities (e.g. approximately 15.4 more private bank branches, 2.3 more cooperative banks, and 2.2 more agricultural credit societies). We also observe improvements in community infrastructure, including 1.1 additional public libraries, 1.4 more public reading rooms, and 0.8 more cinema halls. These are sizable improvements, given that settlements just below the threshold typically have only a handful of such facilities. Results are robust to bandwidth choices and polynomial specifications [^1] Available in forthcoming drafts., and they remain consistent when clustering standard errors at the district or state level. Furthermore, robustness checks support the validity of the design – we find no evidence for manipulation around the thresholds, detect overall balance in baseline characteristics and find null reduced form effects via the direct relationship between crossing the thresholds and outcomes absent of the change in statutory town status.
@@
Our work builds on a growing literature that examines how local institutional capacity and governance shape development outcomes in low- and middle-income countries. besley2010state emphasize that investments in fiscal and legal capacity are key determinants of long-run development, and that variations in local administrative effectiveness can shape the quality of service delivery. Relatedly, gadenne2017tax shows that greater fiscal autonomy can improve accountability and service delivery when citizens can observe how funds are used, while burgess2015value highlights how political incentives affect the allocation of public goods across regions. We contribute to this literature by providing causal evidence on a different but complementary margin — the formal transition of villages to urban local governance. Rather than focusing on local fiscal resources, we study how the administrative act of granting municipal status itself changes public goods provision. By leveraging the multi-threshold eligibility criteria for Census Town classification as a quasi-experiment, we isolate exogenous variation in the likelihood of urban local governance and estimate its local average treatment effect on infrastructure outcomes. This approach provides new evidence on how the extension of urban local governance institutions affects provision of local public goods in urbanizing economies.
@@
The paper contributes to the literature via two main fronts. First, it adds to the causal evidence literature by quantifying the effect of urban governance on local public goods provision in large developing economies. Second, it uses a novel instrument and builds a framework for using multi-dimensional cutoffs in a fuzzy RD design setting, which can be applied to other contexts where policy eligibility and implementation is affected by multiple criteria, allowing researchers to exploit quasi-random variation in policy exposure for causal inference.
Section 2 covers the literature review. Section 3 provides institutional background and describes data construction. Section 4 develops the identification strategy and estimation. Section 5 presents results and robustness. The conclusion discusses implications for urban policy and future research avenues.
2 Literature Review
This section situates our contribution within several strands of literature on local governance, urbanization, and econometric methods.
Urban governance and decentralization.
A large body of work examines whether shifting authority to local governments leads to better public goods provision and welfare outcomes. Classic fiscal federalism theory argues that local governments have informational advantages and can be more responsive to citizen needs (Oates1972). Greater proximity and accountability at the local level may improve allocation of resources to match local preferences. However, other theories warn of potential pitfalls: local elites might capture resources, or small governments may lack administrative capacity and financial resources (BardhanMookherjee2006). Empirical findings on decentralization are mixed. For example, faguet2014can finds that decentralization in Bolivia increased investment in social infrastructure, but similar reforms in other contexts have yielded varying results depending on community participation and oversight. In India, the 73rd and 74th Constitutional Amendments devolved powers to rural panchayats and urban municipalities, respectively, spurring research on their impacts. Studies of village council (panchayat) reforms have found that who leads the council can matter for public goods: for instance, ChattopadhyayDuflo2004 show that mandating women leaders in Indian villages altered the composition of public goods (e.g., more drinking water infrastructure), and (BesleyPandeRao2005) document that more accountable local governments can better target beneficiaries. Still, there is limited causal evidence isolating the effect of having an urban local government in particular – a gap this paper begins to fill.
India’s urban classification and Census Towns.
Our study also relates to research on India’s unique two-tier system of urban classification. Indian settlements can be recognized as urban in two ways: (1) by statutory notification as a municipality or other ULB (a political process controlled by state governments), or (2) by satisfying the Census definition of a “Census Town” based on demographic and economic criteria (a statistical designation by the Census of India). A rich descriptive literature documents the tension between rapid urban population growth and slow municipal incorporation. For example, pradhan2017 highlights the phenomenon of new Census Towns in the 2011 Census – over 3,500 places that became urban by Census criteria since 2001 – many of which were not granted municipal status (“unacknowledged urbanisation”). Kundu2011 and others have argued that state governments often hesitate to create new municipalities due to fiscal and political constraints, leading to a backlog of urbanizing villages that lack urban governance. Our work uses the Census Town thresholds as a natural experiment to study this institutional friction. In doing so, we build on government documentation of these criteria (Registrar General of India, 2011) and recent analyses of the socio-economic differences between statutory towns and Census Towns. By focusing on settlements around the threshold cutoffs, we effectively compare places that are very similar in population and economy – one just meeting the “urban” criteria and one just below – to see if being officially recognized as a town yields different outcomes.
Urban public goods and service delivery.
Our outcome measures contribute to a literature on public service provision in urban areas. Much prior research on urban services in developing countries has relied on household surveys or city-level aggregates. For instance, health and education outcomes are often studied via household data (e.g., the National Family Health Survey or NFHS in India, which provides indicators like immunization rates or school attendance) or via municipal expenditures on services. These approaches, however, can miss granular differences in infrastructure availability at the community level. We instead use the village/town administrative directories to directly measure the presence and quantity of infrastructure (schools, hospitals, banks, libraries, sports facilities, etc.) in each settlement. This complements studies that have examined urban service delivery through other lenses – such as comparisons of slum versus non-slum neighborhoods Auerbach2018 or analyses of the effect of local political incentives on service quality Banerjee2024. By examining a broad set of infrastructure outcomes, our paper provides a holistic view of how municipal status might change the provision of public goods. The evidence on whether municipal governance improves services is limited; for example, some case studies suggest that cities with elected mayors perform better in solid waste management or water provision, but confounding factors abound. Our contribution is to use a causal design to measure these effects at the settlement level.
In summary, our study provides new evidence on the development impacts of urban local governance, leveraging a policy threshold in India’s urban classification system. It bridges the decentralization literature and urbanization debates, introduces a novel multi-threshold RD approach, and contributes a unified data framework for analyzing settlements. Next, we describe the institutional setting and data in more detail.
@@ 3.1 Institutional Background
India’s urban classification rests on two pillars. First, are settlements officially designated as urban by state governments via a notification under municipal law. These include municipal corporations, municipalities, and nagar panchayats222a type of town council for smaller urban areas, and are each known as Statutory Towns (STs). Second, Census Towns (CTs) are settlements that satisfy Census thresholds: total population at least 5,000; population density of at least 400 persons per km²; and at least 75 percent of the male main workforce in non-agricultural activities. CTs are urban for statistical purposes but may continue under rural governance unless notified as STs by the state governments. Crossing the CT thresholds increases the chance of statutory recognition but does not guarantee it as states differ in municipalization policy and timing. Some states have proactive policies to incorporate new urbanizing areas, while others delay or avoid creating new ULBs due to budgetary or political considerations.
@@ 3.2 Local Governance in India
3.2.1 How are Urban Local Bodies (ULBs) formed?
The Nagarpalika Act, which was the 74th Amendment to the Indian Constitution enacted in 1992, provides the legal framework for urban local governance in India (MHA_74thAmendment). This constitutional amendment is supported by state-level municipal acts, such as GoaMunicipalities1968; GujaratMunicipalities1963; KarnatakaMunicipalities1964; MaharashtraMunicipalities1965; TelanganaMunicipalities2019 etc., that outline the specific procedures for forming Urban Local Bodies (ULBs). Although the process differs slightly across states, it typically has similar considerations and involves several steps:
  1. 1.
    Identification and Survey: State governments or local development authorities identify settlements that have urbanized characteristics or meet certain census or urban-classification criteria333as per state’s discretion. (population, density, non-agricultural workforce). A survey may be conducted to assess the settlement’s infrastructure, population, and economic activities.
  2. 2.
    Proposal and Feasibility Study: In many states, the district administration or relevant state department may prepare a proposal for granting municipal status. This includes a feasibility assessment of the settlement’s capacity to sustain urban governance, including revenue potential, administrative requirements, and infrastructure needs.
  3. 3.
    Public Consultation: The proposal may be opened for public consultation where residents and stakeholders can provide feedback. This step also varies by state and may include public hearings or written submissions.
  4. 4.
    State Government Approval: The proposal is submitted to the state government, usually through the Urban Development or Municipal Affairs Department, for review. The state government evaluates political, administrative, and fiscal considerations.
  5. 5.
    Official Notification: Upon approval, the state government issues an official gazette notification declaring the settlement as a statutory town (municipality, municipal corporation, or nagar panchayat, depending on size and status). This notification legally establishes the ULB and defines its boundaries.
  6. 6.
    Election and Formation: Following notification, elections are conducted to form the urban local body. Elected representatives constitute the municipal council or corporation, with a mayor or chairperson as the executive head.
  7. 7.
    Transfer of Functions and Resources: The newly formed ULB assumes responsibility for urban functions as per the 74th Amendment and state municipal acts. For many states, this includes transfer of staff, assets, and revenue sources from previous administrative authorities.
The entire process, from initial identification to operational ULB, can take several years and is subject to state-specific variations in policy, political will, and administrative capacity. Some states proactively municipalize urbanizing settlements, while others delay recognition due to fiscal or political constraints, leading to the phenomenon of Census Towns—settlements that are functionally urban but remain under rural governance.
@@ 3.2.2 What comes with ULB recognition?
When a settlement receives statutory recognition as an Urban Local Body (ULB), it undergoes a fundamental transformation in governance structure, fiscal arrangements, and administrative responsibilities. The 74th Constitutional Amendment Act of 1992 mandates that ULBs assume control over 18 functions listed in the Twelfth Schedule, including urban planning, regulation of land use, roads and bridges, water supply, public health and sanitation, fire services, urban poverty alleviation, and provision of urban amenities (MHA_74thAmendment).
Crucially, ULB recognition brings changes in three key dimensions:
@@
Service Delivery Mandates: Upon municipalization, ULBs inherit responsibility for providing and maintaining urban infrastructure and services. This includes establishing and operating primary schools, health dispensaries, water supply systems, sewerage and drainage networks, street lighting, solid waste management, and fire protection services. State governments typically transfer relevant staff, assets, and liabilities from predecessor rural bodies to the newly formed ULB, though the comprehensiveness and timeliness of such transfers vary considerably across states.
However, the practical impact of ULB recognition is heterogeneous. Well-functioning municipalities in states with strong urban governance traditions may deliver substantially better services, while newly formed ULBs in resource-constrained states may struggle with capacity limitations, inadequate transfers, and weak revenue collection. Our empirical analysis examines whether, on average, statutory recognition translates into measurable improvements in local public goods provision.

### v2

@@
Does Local Urban Governance Status Matter? Evidence from India
Saannidhya Rawat
March 2026
We exploit quasi-random variation around the multi-threshold criteria used to classify Census Towns (CTs) and focus on settlements near the thresholds that are likely to obtain statutory recognition. Using a local fuzzy regression discontinuity design and a multi-threshold criteria, we show that meeting the CT eligibility in 2001 raises the probability of statutory recognition by 2011. Instrumenting statutory recognition with CT eligibility, we estimate the effects of ULB status on local public goods provision: government schools increase by 13.86 (primary), 7.72 (middle), and 4.89 (secondary) units, healthcare infrastructure expands by 2.53 hospitals and 3.00 family welfare centers, and financial access deepens with 4.09 cooperative banks and 2.84 agricultural credit societies. Community amenities also improve, while sports infrastructure declines by 5.71 facilities, consistent with reallocation of urban land. The corresponding reduced-form estimates are directionally consistent and indicate that crossing the CT eligibility frontier improves public goods provision. Our findings indicate that timely municipalization of emerging urban areas can expand provision of public goods.
@@
Corresponding author: Saannidhya Rawat, Department of Economics, Carl H. Lindner College of Business, University of Cincinnati. Email: rawatsa@mail.uc.edu.
@@ 1 Introduction
Urban local bodies (ULBs) are responsible for a wide range of public goods in developing countries such as schools, primary health facilities, sanitation, streets, lighting, and local regulation. Accordingly, whether and how local urban governance status matters for local economic development is a long-standing question, with implications for many rapidly urbanizing economies (Brollo et al., 2013; Faguet, 2012; Patrick and Mothorpe, 2017). However, a key empirical challenge is that governance structures are endogenous: places that are bigger and more developed are more likely to be recognized as urban towns and receive municipal status, while smaller or less developed settlements may not receive such recognition and therefore stagnate even further by being outside the urban system (Gadenne, 2017). As a result, simple cross-sectional comparisons in such contexts will conflate pre-existing differences with the effects of policy. Put differently, governance status is chosen rather than randomly assigned, making causal inference difficult. To address this challenge, researchers have often relied on varying identification strategies. For instance, Lee (2021) finds a positive effect of urban governance on access to water facilities in India by developing a satellite-based measure of confounders such as urbanization and also controlling for population density and other demographic variables in an OLS framework with fixed effects. By contrast, Mukhopadhyay (2017) finds no significant effect of governance status on public goods, such as water and sanitation, in India by comparing the provision of public goods across different types of urban settlements as categorized by their governance status. Meanwhile, although most papers study the impact on the extensive margin, more recently, Narasimhan and Weaver (2024) focus on the intensive margin of public goods provision, examining how governance status may have differential effects on public goods provision depending on the size of the settlement. More broadly, other studies have also explored this question in different contexts and using various methods (Mukhopadhyay, 2015; Mitra and Nagar, 2018), yet the empirical evidence remains mixed.
@@
This paper leverages a sharp institutional feature in India to address this identification problem: the Census Town (CT) classification thresholds that determine eligibility for an area to be considered urban. The Indian Census defines Census Towns as settlements meeting all three criteria: (i) population of at least 5,000 (ii) population density of at least 400 per square kilometer and (iii) at least 75% of the male main workforce in non-agricultural activities. Crossing these multi-dimensional cutoffs sharply increases the probability that a settlement will be formally recognized by state governments as a Statutory Town (ST), i.e. granted a municipality or other ULB status, but notably, it does not guarantee it. We exploit this quasi-random variation in recognition likelihood around the thresholds to construct a local fuzzy regression discontinuity (RD) design. In effect, meeting the Census Town criteria in 2001 serves as an instrument for obtaining statutory recognition in 2011, allowing us to isolate the effect of local urban governance status on public goods offered by the settlements.
@@
The core idea is a fuzzy multi-threshold RD with multiple running variables. Using rich settlement-level microdata from the Socioeconomic High-resolution Rural-Urban Geographic (SHRUG) platform (Asher et al., 2019), we create a running variable called frontier distance using the three CT cutoffs and restrict attention to settlements near these thresholds. We first confirm a strong first-stage relationship: becoming marginally eligible as a Census Town in 2001 leads to a jump in the probability of statutory recognition in 2011. We then estimate the effect of statutory recognition, i.e., ULB status, on a broad set of development indicators measured by the 2011 Indian Census. Our outcome data are assembled by harmonizing the 2011 Village Directory (VD) and Town Directory (TD), which enumerate local infrastructure and amenities. We link these to demographic data from the 2011 Primary Census Abstract (PCA). This provides a comprehensive settlement-level panel of outcomes including educational facilities, health infrastructure, financial institutions, and other community amenities.
@@
Preview of results. We find that crossing the urban eligibility thresholds leads to a pronounced increase in the likelihood of statutory recognition. Meeting all three Census Town criteria in 2001 raises the probability of obtaining statutory recognition by 7.1 percentage points in our preferred local specification, with a first-stage F-statistic of 18.05. Using CT eligibility as an instrument for statutory town status, we find large local average treatment effects on public goods provision: ULB status increases the number of government primary schools by 13.86 per settlement, middle schools by 7.72, and secondary schools by 4.89. We also find positive effects on health and financial infrastructure: statutory recognition increases hospitals by 2.53, family welfare centers by 3.00, cooperative banks by 4.09, and agricultural credit societies by 2.84. Community amenities also improve, with positive estimates for public libraries and reading rooms, while sports infrastructure declines by 5.71 facilities, consistent with land reallocation in formalizing settlements [^1] Due to the small magnitude of the first-stage coefficient, the IV estimates scale reduced-form effects by a relatively large factor. The corresponding reduced-form estimates are directionally consistent and reported in the appendix.. Results are robust to bandwidth choices, and robustness checks support the validity of the design. We find no evidence of manipulation around the thresholds and detect balance in baseline characteristics [^2] Our final regressions control for any remaining differences in these variables..
@@
Our work builds on a growing literature that examines how local institutional capacity and governance shape development outcomes in low- and middle-income countries. Besley and Persson (2010) emphasize that investments in fiscal and legal capacity are key determinants of long-run development, and that variations in local administrative effectiveness can shape the quality of service delivery. Relatedly, Gadenne (2017) shows that greater fiscal autonomy can improve accountability and service delivery when citizens can observe how funds are used, while Burgess et al. (2015) highlights how political incentives affect the allocation of public goods across regions. We contribute to this literature by providing causal evidence on a different but complementary margin: the formal transition of villages to local urban governance status. Rather than focusing on local fiscal resources, we study how the administrative act of granting municipal status itself changes public goods provision. By leveraging the multi-threshold eligibility criteria for Census Town classification as a quasi-experiment, we isolate exogenous variation in the likelihood of statutory recognition and estimate its local average treatment effect on infrastructure outcomes. This approach provides new evidence on how the extension of local urban governance status affects provision of local public goods in urbanizing economies.
@@
This paper contributes to the literature via two main fronts. First, it adds to the literature by quantifying the effect of local urban governance status on local public goods provision in large developing economies. Although Denis and Marius-Gnanou (2010), Pradhan (2013) and Tewari (2020) have noted that many Census Towns remain under rural governance despite meeting Census’s urban criteria, our study is the first to estimate the loss in public goods provision for these areas that are functionally urban but rurally governed. Second, by leveraging the multi-threshold criteria for Census Town classification, our paper uses a novel instrument and highlights the potential for using multi-dimensional cutoffs in a fuzzy RD design setting. This approach can be applied to other contexts where policy eligibility is determined by multiple threshold-based criteria, allowing researchers to exploit quasi-random variation in policy exposure.
@@ 2.1 Institutional Background
India’s urban classification rests on two pillars. The first consists of settlements officially designated as urban by state governments via notification under municipal law. These include municipal corporations, municipalities, and nagar panchayats [^3] A type of town council for smaller urban areas., and are each known as Statutory Towns (STs). The second consists of Census Towns (CTs), which are settlements that satisfy Census thresholds: total population at least 5,000; population density of at least 400 persons per km2; and at least 75 percent of the male main workforce in non-agricultural activities. CTs are urban for statistical purposes but may continue under rural governance unless notified as STs by the state governments. Crossing the CT thresholds increases the chance of statutory recognition but does not guarantee it as states differ in municipalization policy and timing. Some states have proactive policies to incorporate new urbanizing areas, while others delay or avoid creating new ULBs due to budgetary or political considerations.
@@ 2.2 Local Governance in India
The Nagarpalika Act, which was the 74th Amendment to the Indian Constitution enacted in 1992, provides the legal framework for urban local governance in India (Ministry of Home Affairs, Government of India, 2024). This constitutional amendment is supported by state-level municipal acts, such as Government of Goa, 1968; Government of Gujarat, 1963; Government of Karnataka, 1964; Government of Maharashtra, 1965; Government of Telangana, 2019 etc., that outline the specific procedures for forming Urban Local Bodies (ULBs). The entire process, from initial identification to operational ULB, can take several years and is subject to state-specific variations in policy, political will, and administrative capacity. Some states proactively municipalize urbanizing settlements, while others delay recognition due to fiscal or political constraints, leading to the phenomenon of Census Towns—settlements that are functionally urban but remain under rural governance.
@@ 2.2.1 What comes with ULB recognition?
When a settlement receives statutory recognition as an Urban Local Body (ULB), it undergoes a fundamental transformation in governance structure, fiscal arrangements, and administrative responsibilities. The 74th Constitutional Amendment Act of 1992 mandates that ULBs assume control over 18 functions listed in the Twelfth Schedule, including urban planning, regulation of land use, roads and bridges, water supply, public health and sanitation, fire services, urban poverty alleviation, and provision of urban amenities (Ministry of Home Affairs, Government of India, 2024). Crucially, ULB recognition brings changes in three key dimensions:
@@
Service Delivery Mandates: Upon municipalization, ULBs inherit responsibility for providing and maintaining urban infrastructure and services. This includes establishing and operating primary schools, health dispensaries, water supply systems, sewerage and drainage networks, street lighting, solid waste management, and fire protection services.
However, the practical impact of ULB recognition is heterogeneous. Well-functioning municipalities in states with strong urban governance traditions may deliver substantially better services, while newly formed ULBs in resource-constrained states may struggle with capacity limitations, inadequate transfers, and weak revenue collection.